An dude on mars is playing golf. He hits his golf ball with an initial velocity of 50.0m/s at an angle of 30 degrees with respect to the horizon. (gravity on mars = 3.7m/s^2)

a, how far horizontally does the golf ball go?

b, what is the maximum height of the golf ball during its flight.

c, how long is the ball in the "air"?
Answer #1

A) We know that the horizontal distance

R=u2 sin(2)/g

here u=50 m/s , g=3.7 m/s2 , theta =30o

R=(50)2 sin(2*30o)/3.7

R=585.15 m

b) We know that

Hmax =u2 sin2()/2g

=502 sin2(30o)/(2*3.7)

Hmax =84.46 m

c) time for which the ball in air

t=2u sin()/g

t=2*50 sin(30o)/3.7

t=13.514 sec.